OpenStudy (anonymous):

4200 ft. apart with height of 546 ft above roadway...cable is 15ft above roadway...I have vertex at (0,15) and (2100, 546) , (-2100, 546) but can't get equation to work

OpenStudy (amistre64):

gonna have to chase the question am i...... stay put! lol

OpenStudy (amistre64):

those are some awfully big number, this could get messy :)

OpenStudy (anonymous):

yes I've worked it out several times

OpenStudy (amistre64):

your values are good... (2100^2)a +(2100)b + c = 546 (-2100^2)a -(2100)b + c = 546 0a +0b + c = 15 we see the c = 15 here right?

OpenStudy (amistre64):

(2100^2)a +(2100)b + c-c = 546 -15 (2100^2)a +(2100)b = 531 (-2100^2)a -(2100)b + c-c = 546-15 (2100^2)a -(2100)b = 531 --------------------------- 531 -2100b a = ----------- (2100)^2 substitute this into the next equation right? (2100^2)a +(2100)b = 531 ------------------------------

OpenStudy (amistre64):

(2100^2)(531 -2100b) -------------------- -(2100)b = 531 (2100)^2 531 -2100b - 2100b = 531 it says b = 0 :)

OpenStudy (amistre64):

which is what you would expect it to be since we are on the y axis already and just raised by 15 right?

OpenStudy (anonymous):

so we are using y-y1=c(x-x1)2 yes, y-intercept is (0, 15)

OpenStudy (amistre64):

531 a = ---------- (2100)^2 \[y = \frac{531}{2100^2} x^2 + 15\]

OpenStudy (amistre64):

see if this fits :)

OpenStudy (amistre64):

when x = 2100 we get: 531 + 15 = 546 right?

OpenStudy (amistre64):

come on..you know im right....tell me i did good :)

OpenStudy (anonymous):

I was squaring 2100 and got lost in math... YES you're right...AWESOME!!!! THANKS ALOT!

OpenStudy (anonymous):

I was comparing my work to yours...all different attempts , lol

OpenStudy (amistre64):

lol...... thanx :) you had it tight to begin with, dont let the numbers scare you ;)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you have time for another